Science Report – 03 May 2018

1. Emergency EVAC EVA: No EVAs occurred today due to visit from an elementary school class.

2. Sunspot and Solar Flare Monitoring: No pictures were taken with the camera today, because we hosted a group of school kids on a tour of the entire campus.

3. Shortwave Texting on EVA’s: No EVAs occurred today due to visit from an elementary school class.

4. GPS Route Measurements: No EVAs occurred today due to visit from an elementary school class.

5. Waterless dish cleaning: Bacteria growth and monitoring will continue over the last two days of the mission. The study is continuing to show promising results regarding the use of UV and a cloth to clean dishes on mars without the use of water.

6. The MAG (Maximum Absorbency Garment) utilization study: The study has been scrapped. Results will be from blue crew only.

7. Spacesuit Helmet Fogging: Study has been discontinued after 7 days of data collection.

8. Crew Comfort, Health and Safety on EVA: No EVAs occurred today due to visit from an elementary school class.